# SCI: An Equilibrium for Signal Intelligence

### v1

@@ B.8 Controller Hyperparameters
$SP^{*}$ (target) 0.95 Target signal-perception score
@@ C.1 Datasets
CHB-MIT EEG PhysioNet, 10.13026/C2K01R Seizure detection Open Data Commons
@@
MIMIC-III Waveform PhysioNet, 10.13026/C2294B Alarm triage Credentialed access
IMS/NASA Bearing NASA PCoE Fault detection, RUL Public domain
PHM Tool Wear PHM Society 2010 Tool wear prediction Academic use
NOAA Climate Indices NOAA CPC Anomaly detection Public domain
IRIS Seismic Data IRIS Earthquake detection IRIS Data Policy
Table 8: Summary of datasets, tasks, and access conditions used in SCI experiments.
@@ C.3 Train–online split
Let $V(t)=\tfrac{1}{2}(\Delta SP(t))^{2}$ with $\Delta SP(t)=SP^{*}(t)-SP(t)$ and the projected update
@@
Assume (A1)–(A5) from §5.4: $L$-smoothness of $SP(\Theta)$, bounded gradients $\|\nabla_{\Theta}SP\|\leq G$, bounded human signal $\|u_{h}(t)\|\leq U$, slowly varying $w_{f}(t)$ and $SP^{*}(t)$. Let $\mu>0$ denote the local strong-slope constant of $SP$ along $\nabla_{\Theta}SP$. By $L$-smoothness and the non-expansiveness of $\mathrm{Proj}_{\mathcal{C}}$,

### v2

@@ B.8 Controller Hyperparameters
$SP^{\star}$ (target) 0.95 Target signal-perception score
@@ C.1 Datasets
MNIST digits LeCun et al. (1998) Digit classification Public
@@
IMS/NASA Bearing NASA PCoE Fault detection Public domain
CHB-MIT EEG PhysioNet, 10.13026/C2K01R Seizure detection (planned) Open Data Commons
MIMIC-III Waveform PhysioNet, 10.13026/C2294B Alarm triage (planned) Credentialed access
PHM Tool Wear PHM Society 2010 Tool wear prediction (planned) Academic use
NOAA Climate Indices NOAA CPC Anomaly detection (planned) Public domain
IRIS Seismic Data IRIS Earthquake detection (planned) IRIS Data Policy
Table 8: Summary of datasets used in the current SCI experiments (top) and additional domains targeted for future evaluation (bottom). The empirical results in §7 are restricted to MNIST, MIT–BIH ECG, and IMS/NASA bearings.
@@ C.3 Train–online split
Let $V(t)=\tfrac{1}{2}(\Delta SP(t))^{2}$ with $\Delta SP(t)=SP^{\star}(t)-SP(t)$ and the projected update
@@
Assume (A1)–(A5) from §5.4: $L$-smoothness of $SP(\Theta)$, bounded gradients $\|\nabla_{\Theta}SP\|\leq G$, bounded human signal $\|u_{h}(t)\|\leq U$, slowly varying $w_{f}(t)$ and $SP^{\star}(t)$. Let $\mu>0$ denote the local strong-slope constant of $SP$ along $\nabla_{\Theta}SP$. By $L$-smoothness and the non-expansiveness of $\mathrm{Proj}_{\mathcal{C}}$,